# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
Fig. 3c depicts $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as function of $\tau_{\mathrm{h}}$ for three different out-of-plane magnetic fields. All three datasets show an exponential decay of the occupation probability as a function of $\tau_{\mathrm{h}}$. From an exponential fit (solid line), we determine for example $T_{1}=4.0~{}\mu$s at $B_{\perp}=0.175$ T. $T_{1}$ decreases with increasing $B_{\perp}$ and reaches a value of 845 ns at $B_{\perp}=0.45$ T. A single charge carrier occupying |$\ket{K\uparrow}$ may relax into a lower lying state either by pure valley relaxation (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\uparrow}$) or by additionally flipping the spin (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\downarrow}$). Relaxation processes requiring a single valley flip are expected to be faster than processes that require both a spin and valley flip. This is supported by the fact that spin relaxation times between hundreds of microseconds to up to 50 ms have been recently reported in BLG  Banszerus et al. (2022); Gächter et al. (2022). Hence, we conclude that $T_{1}$ extracted from Fig. 3c must be limited by the valley relaxation time. In Fig. 4, we plot the valley relaxation time $T_{1}$ (data points) extracted from exponential fits, as exemplarily shown in Fig. 3c as a function of $B_{\perp}$ and as a function of the energy splitting $\Delta E_{\mathrm{v}}$. When decreasing $B_{\perp}$ from 0.7 T to about 0.15 T, $T_{1}$ increases from below 0.5 $\mu$s to about 7 $\mu$s, while at even lower $B_{\perp}$, the relaxation rate decreases again to $T_{1}\sim$$\mu$s at 80 mT (see gray data points).
@@
To gain a better understanding of the experimental $T_{1}$ results, we compare them with theory and therefore model the system by the Hamiltonian $H=H_{\rm QD}+H_{\rm EPC}+H_{KK^{\prime}}$, where $H_{\rm QD}$ describes a single electron or hole in the BLG QD and $H_{\rm EPC}=\sum_{\lambda q}H_{\rm EPC}^{\lambda q}$ the electron-phonon coupling. Furthermore, we allow a mixing between the two valleys described by the intervalley coupling term $H_{KK^{\prime}}=\Delta_{KK^{\prime}}\tau_{x}/2$ where $\tau_{x}$ is the Pauli matrix acting on the valley degree of freedom. For simplicity, the electrostatic confinement is modeled by a finite circularly symmetric step potential $U({\bf r})$ with potential depth $U_{0}\approx 39.6$meV and QD radius $r=25$nm. This yields a valley g-factor of $g_{\text{v}}=30$, which is in good agreement with the experimental data (c.f. Fig. 2e). We consider transitions between states with equal spin, but opposite valley degree of freedom mediated by coupling to in-plane acoustic phonons arising either from the deformation potential (characterized by the coupling strength $g_{1}$) or from bond-length change (coupling strength $g_{2}$ohier et al. (2014). The Hamiltonian describing coupling to phonons in the mode $\lambda$ with wave vector $q$ has the form $H_{\rm EPC}^{\lambda q}=c_{q}(g_{1}a_{1}\sigma_{0}+g_{2}a^{\prime}_{2}\sigma_{x}+g_{2}a^{\prime\prime}_{2}\sigma_{y})(e^{i{\bf q}\cdot{\bf r}}b^{\dagger}_{\lambda q}-e^{-i{\bf q}\cdot{\bf r}}b_{\lambda q})$ with $\sigma_{x,y,z}$ the Pauli matrices for the sublattice degree of freedom Ando (2005); Mariani and von Oppen (2008), and $c_{q}=\sqrt{q/A\rho v_{\rm\lambda}}$, with $A$ the area of the BLG sheet, $\rho$ the mass density of BLG, and $v_{\rm\lambda}$ the sound velocity; $a_{1,2}$ are phase factors and $b_{\lambda q}$ and $b^{\dagger}_{\lambda q}$ are the phonon ladder operators Wang and Burkard . Using Fermi’s golden rule, we calculate the valley relaxation times $T_{1}$ between initial and final eigenstates |$\ket{i}$ and |$\ket{f}$ of the Hamiltonian $H_{\rm QD}+H_{KK^{\prime}}$ with opposite valley quantum number and eigenenergies $\varepsilon_{i}$ and $\varepsilon_{f}$,
$\frac{1}{T_{1}}=2\pi A\sum_{\lambda}\int\frac{d^{2}q}{(2\pi)^{2}}|\bra{i}H_{\mathrm{EPC}}^{\lambda q}\ket{f}|^{2}\delta(\varepsilon_{f}-\varepsilon_{i}+v_{\mathrm{\lambda}}q).$ (2)
@@
We only take into account the emission of phonons (with energy $v_{\mathrm{\lambda}}q$) as the thermal energy is significantly smaller than the valley splitting. To quantify the electron-phonon coupling strength, we perform the least square fit to the experimental data using $g_{1}$ and $g_{2}$ as free fit parameters. Our model is in good qualitative and quantitative agreement with the data taken above $B_{\perp}=0.1$ T, where increasing $B_{\perp}$ results in decreasing $T_{1}$, while it cannot explain the decrease in $T_{1}$ observed for $B_{\perp}<0.1~{}$T, suggesting that other mechanisms dominate $T_{1}$ in this regime. We speculate that the discrepancy between our model and the data in this regime may arise from a hot spot Yang et al. (2013) or charge noise Huang and Hu (2014); Hosseinkhani and Burkard (2021). Consequently, we have restricted the fit to the data taken for $B_{\perp}\geq 0.1$ T. The fit yields coupling parameters of $g_{1}~{}=~{}50~{}$eV, $g_{2}=5.4~{}$eV. It is noteworthy that both parameters are in good agreement with the reported literature values, which include values in the range of $20$ to $50$ eV for $g_{1}$ Park et al. (2014); Mariani and von Oppen (2010); Chen et al. (2008); Hwang and Das Sarma (2008); Suzuura and Ando (2002) and values in the range of $1.5$ to $5$ eV for $g_{2}$ Suzuura and Ando (2002); Sohier et al. (2014), where the wide range of values is partly due to the dependence of the deformation potential on screening and doping Park et al. (2014); Mariani and von Oppen (2010). The black solid line in Fig. 4 corresponds to the contributions from both, the deformation potential coupling and from bond length change coupling, while the blue and red dashed lines show the individual contributions, respectively (see labels in Fig. 4). In the calculation, the intervalley coupling, which is mainly responsible for the absolute $T_{1}$ values but does not enter the functional $B_{\perp}$ dependence, was set to $\Delta_{KK^{\prime}}=50~{}\mu$eV.
@@
At larger magnetic fields, $B_{\perp}\gtrsim 0.5~{}$T, $T_{1}$ is predominantly limited by electron-phonon coupling via the deformation potential, whereas at smaller fields it is limited by the coupling due to bond length change. This transition occurs due to the crossover between the dipole and the higher multipole regimes for the bond-length change coupling if the phonon wavelength $\lambda\approx 2\pi\hbar v_{\rm\lambda}/(g_{\rm v}\mu_{B}B_{\perp})$ is comparable to the QD radius $r$, where $qr=2\pi r/\lambda\approx 1$. Hence, the crossover occurs around the critical field $B^{c}_{\perp}\approx\hbar v_{\rm\lambda}/(g_{\rm v}\mu_{\mathrm{B}}r)\approx$ 0.3 T. The assumed QD radius of $r=25$ nm is in agreement with the lithographic device dimensions as well as with the confinement size giving rise to a valley g-factor of $g_{\text{v}}=30$ in excellent agreement with experiment. The gray horizontal bar in Fig. 4 depicts the range of $B^{c}_{\perp}$ assuming the estimate of $r$ to deviate by a factor of 2 highlighting that the transition region is well within the experimentally investigated $B$- field range.
@@
The long single-particle valley lifetimes in BLG QDs of up to 7 $\mu$s demonstrated by our work confirms that the valley degree of freedom is indeed an interesting candidate for implementing qubits. This potential is furthermore underlined by a recent experiment showing long relaxation times from valley triplet to valley singlet states Garreis et al. (2024). By fitting to the experimental data, we confirm that over a wide magnetic field range, electron-phonon coupling mediated by bond length change and the deformation potential limit the relaxation time. As the valley magnetic moment is typically one to two orders of magnitude larger than the magnetic moment associated with the electron spin, we anticipate gate operation times of a valley qubit to be much faster than those of a spin qubit, potentially compensating for the shorter relaxation times. The magnitude of the valley magnetic moment can be adjusted all-electrically, which could provide a way to realize control over a single valley without the need for microwave bursts, micromagnets or ESR strips, and enable well-controlled qubit addressability. A crucial follow-up experiment is the determination of coherence times ($T_{2}^{*}$ and $T_{2}$) in this system, potentially using a recently demonstrated particle electron-hole blockade as a readout method Banszerus et al. (2023).
Methods
[…]
In order to perform RF gate modulation, the sample is mounted on a home-built printed circuit board (PCB). All DC lines are low-pass-filtered (10 nF capacitors to ground). All FGs are connected to on-board bias-tees, allowing for AC and DC control on the same gate (see Fig. 2a). All AC lines are equipped with cryogenic attenuators of -26 dB. $V_{\text{AC}}$ refers to the AC voltage applied prior to attenuation. The measurements are performed in a ${}^{3}$e/${}^{4}$e dilution refrigerator at a base temperature of approximately 20 mK and at an electron temperature of around 60 mK. The current through the device is amplified and converted into a voltage with a home-built I-V converter at a gain of $10^{8}$. Throughout the experiment, a constant back gate voltage of $V_{\mathrm{BG}}=5.025~{}$V and a split gate voltage of $V_{\mathrm{SG}}=-5.435~{}$V is applied to define a n-type channel between source and drain. The four gates acting as barrier gates to the QD (see Fig. 2a, yellow and blue color coding) are biased with $-6.05\pm 0.1$ V, -4.95 V, -5.18 V and $-6.15\pm 0.1$ V, respectively. The voltages are adjusted to compensate for the influence of $B_{\perp}$ on the tunnel coupling.
The Hamiltonian $H_{\mathrm{BLG}}$ describing the band structure of bulk BLG is given by
$H_{\mathrm{BLG}}(\mathbf{k})=\begin{bmatrix}\Delta&\gamma_{0}p&\gamma_{4}p*&\gamma_{1}\\ \gamma_{0}p*&\Delta&\gamma_{3}p&\gamma_{4}p*\\ \gamma_{4}p&\gamma_{3}p*&-\Delta&\gamma_{0}p\\ \gamma_{1}&\gamma_{4}p&\gamma_{0}p*&-\Delta\end{bmatrix}$ (3)
with the displacement field $2\Delta$ and the hopping parameters $\gamma_{0}=2.6$ eV, $\gamma_{1}=0.339$ eV, $\gamma_{3}=0.28$ eV and $\gamma_{4}=-0.14$ eV. The momentum $p(\mathbf{k})=-\sqrt{3}a(\tau k_{x}-ik_{y}-ixB_{\perp}e/2-\tau yB_{\perp}e/2)/2$ with the valley index $\tau=\pm 1$ and the lattice constant $a=2.46$ includes the valley Zeeman effect McCann and Koshino (2013); Konschuh et al. (2012); Wang and Wu (2013).
@@ Acknowledgements
The authors thank A. Hosseinkhani for fruitful discussions and F. Lentz, S. Trellenkamp, M. Otto and D. Neumaier for help with sample fabrication. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 881603 (Graphene Flagship) and from the European Research Council (ERC) under grant agreement No. 820254, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, through DFG (STA 1146/11-1), and by the Helmholtz Nano Facility Albrecht et al. (2017). K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 20H00354 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. L.W. and G.B. acknowledge support from DFG Project No. 425217212, SFB 1432.
@@
[…]

### v2

@@ Abstract
Fig. 3(c) depicts $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as function of $\tau_{\mathrm{h}}$ for three different out-of-plane magnetic fields. The datasets show an exponential decay of the occupation probability as a function of $\tau_{\mathrm{h}}$. An exponential fit (solid line), yields for example $T_{1}=4.0~{}\mu$s at $B_{\perp}=0.175$ T. $T_{1}$ decreases with increasing $B_{\perp}$ and reaches a value of 845 ns at $B_{\perp}=0.45$ T. A single charge carrier occupying |$\ket{K\uparrow}$ may relax into a lower lying state either by pure valley relaxation (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\uparrow}$) or by additionally flipping the spin (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\downarrow}$). Relaxation processes requiring a single valley flip are expected to be faster than processes that require both a spin and valley flip. This is supported by recently reported spin relaxation times between hundreds of microseconds to up to 50 ms, for energy splittings, $\Delta E_{\mathrm{S}}>200$ $\mu$eV [33, 34]. For the pure spin relaxation between |$\ket{K^{\prime}\downarrow}$ and the GS, no relaxation could be observed over the whole range of investigated $\tau_{\mathrm{h}}$ and $B_{\perp}$ (the amplitude of |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$ is constant as a function of $\tau_{\mathrm{h}}$, see Supplementary Fig. S1). Hence, we conclude that $T_{1}$ extracted from Fig. 3(c) must be limited by the valley relaxation time. In Fig. 4, we plot the valley relaxation time $T_{1}$ extracted from exponential fits, as exemplarily shown in Fig. 3(c) as a function of $B_{\perp}$ and as a function of the energy splitting $\Delta E_{\mathrm{v}}$. When decreasing $B_{\perp}$ from 0.7 T to about 0.15 T, $T_{1}$ increases from below 0.5 $\mu$s to about 7 $\mu$s, while at even lower $B_{\perp}$, the relaxation rate decreases again to $T_{1}\sim$$\mu$s at 80 mT (gray).
@@
To gain a better understanding of the experimental results, we compare them with theory. We model the system by the Hamiltonian $H=H_{\rm QD}+H_{\rm EPC}+H_{KK^{\prime}}$, where $H_{\rm QD}$ describes a single electron or hole in the BLG QD and $H_{\rm EPC}=\sum_{\lambda q}H_{\rm EPC}^{\lambda q}$ the electron-phonon coupling. Furthermore, we allow a mixing between the two valleys described by the intervalley coupling term, $H_{KK^{\prime}}=\Delta_{KK^{\prime}}\tau_{x}/2$ with the Pauli matrix $\tau_{x}$ acting on the valley degree of freedom. For simplicity, we model electrostatic confinement by a finite circularly symmetric step potential $U({\bf r})$ with potential depth $U_{0}\approx 39.6$meV and $r=25~{}$nm. This yields a valley g-factor of $g_{\text{v}}=30$, in good agreement with the experiment (c.f. Fig. 2(e)). We consider transitions between states with equal spin, but opposite valley degree of freedom mediated by coupling to in-plane acoustic phonons arising either from the deformation potential (coupling strength $g_{1}$) or from bond-length change ($g_{2}$[36]. Since we operate in the low energy limit, we only consider acoustic phonons, while out-of-plane acoustic (ZA) phonons are supposed to be quenched in supported, especially in encapsulated graphene [37, 25]. The Hamiltonian describing coupling to phonons in the mode $\lambda$ with wave vector $q$ has the form $H_{\rm EPC}^{\lambda q}=c_{q}(g_{1}a_{1}\sigma_{0}+g_{2}a^{\prime}_{2}\sigma_{x}+g_{2}a^{\prime\prime}_{2}\sigma_{y})(e^{i{\bf q}\cdot{\bf r}}b^{\dagger}_{\lambda q}-e^{-i{\bf q}\cdot{\bf r}}b_{\lambda q})$ with $\sigma_{x,y,z}$ the Pauli matrices for the sublattice degree of freedom [38, 39], and $c_{q}=\sqrt{q/A\rho v_{\rm\lambda}}$, with $A$ the area of the BLG sheet, $\rho$ the mass density of BLG, and $v_{\rm\lambda}$ the sound velocity; $a_{1,2}$ are phase factors and $b_{\lambda q}$ and $b^{\dagger}_{\lambda q}$ are the phonon ladder operators [25]. Using Fermi’s golden rule, we calculate the valley relaxation times $T_{1}$ between initial and final eigenstates |$\ket{i}$ and |$\ket{f}$ of the Hamiltonian $H_{\rm QD}+H_{KK^{\prime}}$ with opposite valley quantum number and eigenenergies $\varepsilon_{i}$ and $\varepsilon_{f}$,
$\frac{1}{T_{1}}=2\pi A\sum_{\lambda}\int\frac{d^{2}q}{(2\pi)^{2}}|\bra{i}H_{\mathrm{EPC}}^{\lambda q}\ket{f}|^{2}\delta(\varepsilon_{f}-\varepsilon_{i}+v_{\mathrm{\lambda}}q).$
@@
We only take into account the emission of phonons (with energy $v_{\mathrm{\lambda}}q$) as the thermal energy is significantly smaller than the valley splitting. To quantify the electron-phonon coupling strength, we perform the least square fit to the experimental data using $g_{1}$ and $g_{2}$ as free fit parameters. Our model is in good qualitative and quantitative agreement with the data taken above $B_{\perp}=0.1$ T, where increasing $B_{\perp}$ results in decreasing $T_{1}$, while it cannot explain the decrease in $T_{1}$ observed for $B_{\perp}<0.1~{}$T, suggesting that other mechanisms dominate in this regime. We speculate that the discrepancy between our model and the data in this regime may be due to a hot spot [40], thermal broadening induced charge noise [41, 42] or $1/f$ charge noise, which is expected to become dominant in the small energy range, as discussed in more detail in ref. [25]. Consequently, we have restricted the fit to the data for $B_{\perp}\geq 0.1$ T, which yields coupling parameters of $g_{1}=50~{}$eV, $g_{2}=5.4~{}$eV. It is noteworthy that both parameters are in good agreement with literature, which includes values in the range of $20$ to $50$ eV for $g_{1}$ [43, 44, 45, 46, 47] and values in the range of $1.5$ to $5$ eV for $g_{2}$ [47, 36], where the wide range of values is partly due to the dependence of the deformation potential on screening and doping [43, 44]. The black solid line in Fig. 4 corresponds to the contributions from both, the deformation potential coupling and from bond length change coupling, while the dashed lines show the individual contributions, respectively (see labels). In the calculation, the intervalley coupling, which is mainly responsible for the absolute values of $T_{1}$ but does not enter the functional $B_{\perp}$ dependence, was set to $\Delta_{KK^{\prime}}=50~{}\mu$eV.
@@
At larger magnetic fields, $B_{\perp}\gtrsim 0.5~{}$T, $T_{1}$ is predominantly limited by electron-phonon coupling via the deformation potential, whereas at smaller fields it is limited by the coupling due to bond length change. This transition occurs due to the crossover between the dipole and the higher multipole regimes for the bond-length change coupling if the phonon wavelength $\lambda\approx 2\pi\hbar v_{\rm\lambda}/(g_{\rm v}\mu_{B}B_{\perp})$ is comparable to the QD radius $r$, where $qr=2\pi r/\lambda\approx 1$. Hence, the crossover occurs around the critical field $B^{c}_{\perp}\approx\hbar v_{\rm\lambda}/(g_{\rm v}\mu_{\mathrm{B}}r)\approx$ 0.3 T. The assumed QD radius of $r=25$ nm is in agreement with the lithographic device dimensions as well as with the confinement size, giving rise to a valley g-factor of $g_{\text{v}}=30$ in excellent agreement with experiment. The gray bar in Fig. 4 depicts the range of $B^{c}_{\perp}$ assuming the estimate of $r$ to deviate by a factor of 2 highlighting that the transition region is well within the experimentally investigated $B_{\perp}$ range. The long single-particle valley lifetimes in BLG QDs of up to 7 $\mu$s make BLG a promising candidate for valleytronic applications and confirm that the valley degree of freedom is indeed interesting for the realization of qubits, where the $T_{1}$ time sets an upper limit on the coherence time $T_{2}^{*}$.
@@
This potential is furthermore underlined by a recent experiment showing long relaxation times from valley triplet to valley singlet states [48]. Fits to the experimental data confirm that electron-phonon coupling mediated by bond length change and the deformation potential limit the relaxation time over a wide magnetic field range. As the valley magnetic moment is typically one to two orders of magnitude larger than the magnetic moment associated with the electron spin, we anticipate gate operation times of a valley qubit to be much faster than those of a spin qubit, potentially compensating for the shorter relaxation times. The magnitude of the valley magnetic moment can be adjusted all-electrically [30], which could provide a way to realize control over a single valley without the need for microwave bursts, micromagnets or ESR strips, and enable well-controlled addressability. Crucial follow-up experiments include the determination of the coherence times ($T_{2}^{*}$ and $T_{2}$) and insights into their limiting mechanisms such as charge noise, as well as the understanding of the upper bound of the recently reported long relaxation times of Kramers states at small energy splittings [49].
@@ Acknowledgements
The authors thank A. Hosseinkhani for fruitful discussions and F. Lentz, S. Trellenkamp, M. Otto and D. Neumaier for help with sample fabrication. This project has received funding from the European Research Council (ERC) under grant agreement No. 820254, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, through DFG (STA 1146/11-1), and by the Helmholtz Nano Facility [50]. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. L.W. and G.B. acknowledge support from DFG Project No. 425217212, SFB 1432.
@@
C.S. designed and directed the project; L.B., K.H., S.M. fabricated the device, L.B., K.H. and C.V. performed the measurements and analyzed the data. K.W. and T.T. synthesized the hBN crystals. L.W. and G.B. performed calculations of the T1 time. G.B., C.V. and C.S. supervised the project. L.B., K.H., L.W., G.B, C.V. and C.S. wrote the manuscript with contributions from all authors. L.B. and K.H. contributed equally to this work. Correspondence should be addressed via e-mail to katrin.hecker@rwth-aachen.de